hello. dollar $ ampersand & procent % matje # leftbrace { rightbrace } underscore

dagger †ddagger ‡section §paragraph ¶copyright \begin{matrix}c \\ ◯ \end{matrix}pound £

• item 1

• – item 2.1

bye.

maths:

1 + 2 = 5/3 ∗ 1 − 1